# Disentangling representations of retinal images with generative models

### v2

@@ Acknowledgements
We thank Holger Heidrich for discussing project directions and implementation details. We also thank Patrick Köhler, Kyra Kadhim, Simon Holdenried-Krafft, and Jeremiah Fadugba for helpful discussions and Kyra Kadhim, Julius Gervelmeyer, and Holger Heidrich for comments on the manuscript. This project was supported by the Hertie Foundation and by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy with the Excellence Cluster 2064 “Machine Learning — New Perspectives for Science”, project number 390727645. This research utilized compute resources at the Tübingen Machine Learning Cloud, INST 37/1057-1 FUGG. PB is a member of the Else Kröner Medical Scientist Kolleg “ClinbrAIn: Artificial Intelligence for Clinical Brain Research”. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting SM.
@@ Appendix A Toy example for disentanglement loss
Although distance correlation is simple to compute, it is able to measure nonlinear dependencies between variables. With a simple toy example we show how distance correlation handles dependencies between subspaces compared to two linear dependency measures. We sampled 1,000 points from two 2-dimensional subspaces $w_{1}=[w_{1,1},w_{1,2}]\in\mathbb{R}^{2}$ and $w_{2}=[w_{2,1},w_{2,2}]\in\mathbb{R}^{2}$ and created dependencies between them (Fig. A.1a, main diagonals in the pair plots). We defined an optimization problem, where points should be moved in order to minimize different dependence measures (Fig. A.1b-d). We optimized point coordinates with PyTorch (Paszke et al., 2019) and an Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.99,\epsilon=10^{-8}$) with a learning rate of 0.05 for 500 steps. When we optimizes for linear dependence measures like Gaussian mutual information (GMI) or collapsed Gaussian mutual information (C-GMI) the dependence measure only decreased for linear dependencies (Fig. A.1b,c, distance correlation values below points). GMI and C-GMI are two mutual information estimators that assume that the subspace vectors $w_{1}$ and $w_{2}$ follow a multivariate normal distribution. For a detailed background on GMI and C-GMI, see the next Sec. A.1. However, when we optimized for distance correlation, the dependence measure also dropped for nonlinear dependencies (Fig. A.1d, distance correlation values below points).
Figure A.1: Toy example for different dependency measures. We sampled 1,000 points from two 2-dimensional subspaces $w_{1}=[w_{1,1},w_{1,2}]\in\mathbb{R}^{2}$ and $w_{2}=[w_{2,1},w_{2,2}]\in\mathbb{R}^{2}$ and created dependencies between the main diagonals in the pair plots of a. We defined an optimization problem where we changed point coordinates to minimize various dependency measures. Below the points we show distance correlation as a dependence measure. In panels b and c we show the displacement of points when minimizing the linear dependency measures GMI and C-GMI, respectively. In panel d we see that, compared to GMI and C-GMI, distance correlation can also effectively disentangle nonlinear dependencies between subspaces.
@@ A.1 Gaussian mutual information
$\text{GMI}(w_{1},w_{2})$ $=\frac{1}{2}\left(\log\left(\frac{\det\Sigma_{1}\det\Sigma_{2}}{\det\Sigma}\right)\right)$
@@
Approximating this measure over batches leads to numerically instabilities for covariance matrix estimation. Hence, we simplify the computation even more by taking the sum of the variables over their feature dimension $\hat{w_{1}}=\Sigma_{i=0}^{d_{1}}w_{1,i}$ and $\hat{w_{2}}=\Sigma_{i=0}^{d_{2}}w_{2,i}$. Since we collapse feature dimensions here, we name this measure collapsed Gaussian mutual information (C-GMI).
@@
Linear dependency measure.
When we assume $w_{1}$ and $w_{2}$ follow a multi-variate normal distribution, the measure GMI measure is only dependent on the covariance $\Sigma$ of the joint distribution $p(w_{1},w_{2})$
@@
Hence, this measure considers only linear dependencies in the uncollapsed and collapsed form.
@@ Appendix E Failure cases
We qualitatively investigated the limitation of our generative model by checking for failure cases of a model trained for age-camera-identity subspace disentanglement. We randomly sampled latent vectors from a standard normal distribution $z\sim\mathcal{N}(0,I)\in\mathcal{R}^{32}$ over different training epochs. We chose our best model in terms of validation loss. At the beginning of the training, when the generative model had not yet converged, the image features were cartoon-like (Fig. E.2a, epoch 9 and 17). In later stages of training, the fundus images got more realistic, however, we discovered some unrealistic features, such as images that were not perfectly circular (Fig. E.2b, first row, left edges), second optic discs (Fig. E.2b, first row, right center), or extreme colors and shapes (Fig. E.2b, second row). Sometimes images were generated with the non-black pixels in the background (Fig. E.2c) and few images even missed important anatomical structures such as the optic disc and the macula (Fig. E.2d). In some cases, we saw fundus images where the optic disc was on the wrong side (Fig. E.2e), and sometimes the images even looked very similar even though they were generated from different latent spaces (Fig. E.2e, first versus second row). We even discovered a few blurry images of poor quality (Fig. E.2f).
@@ Appendix F Different image resolutions
We trained generative models with age-camera-identity disentanglement with different image resolutions and compared their performance with several metrics: disentanglement measure as the mean of subspace distance correlation values, age and camera encoding, both measured as the above chance level of accuracy, and the FID score as a metric of image quality. Here, we could see that our disentanglement measure was similar over different image resolutions (Fig. F.3, a), which was not surprising, since we used the same 32-dimensional latent space for each resolution. Age- and camera-encoding, on the other hand, improves for higher resolution (Fig. F.3, a). One possible explanation could be that it is easier to extract age and camera features from larger images. Finally, the FID score became worse (higher) with increasing resolution (Fig. F.3, a). Better (lower) FID scores tend to show greater sample diversity, however, the individual images were of better quality at higher resolution (Fig. F.3, b). Also, as mentioned above, we used the same 32-dimensional latent space for each image resolution. For a better FID score at higher image resolutions, larger latent spaces may be required.
@@ Appendix G Multiple subspaces
To show that our model can handle combinations of multiple attributes, we trained generative models with subspaces for age, ethnicity, and camera encoding simultaneously. Therefore, compared to our previous experiments with 3 subspaces, we scaled up to 4 subspaces, resulting in 6 subspace combinations for distance correlation minimization (Eq. 4). Here we could see that despite training on 4 subspaces, our model still showed comparable performance: the kNN classifier performances on off-diagonals in our confusion matrix of label-subspace combinations for a distance correlation minimization model dropped similarly as in the experiments with only 3 subspaces (Fig. G.4, a versus b). In addition, the values of the distance correlation between the individual subspace combinations fell from an average of 0.42 to 0.28, without major outliers (Fig. G.4, b).
Figure G.4: Disentanglement performance of a generative model with multiple subspaces. We trained generative models on retinal images, featuring four subspaces: $w_{\text{age}}$ encoding the age class $y_{\text{age}}$, $w_{\text{ethn}}$ encoding patient ethnicity $y_{\text{ethn}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{id}}$ serving as a patient-identity subspace. In a and b, we report kNN classifier accuracy improvement over chance level accuracy. The baseline model in a involved training linear classifiers on the first three subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces for two model configurations.

### v3

@@ Acknowledgements
We thank Francesco Locatello for discussing details about the background on causal representation learning and we thank Holger Heidrich for discussing project directions and implementation details. We also thank Patrick Köhler, Kyra Kadhim, Simon Holdenried-Krafft, and Jeremiah Fadugba for helpful discussions and Kyra Kadhim, Julius Gervelmeyer, and Holger Heidrich for comments on the manuscript. This project was supported by the Hertie Foundation and by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy with the Excellence Cluster 2064 “Machine Learning — New Perspectives for Science”, project number 390727645. This research utilized compute resources at the Tübingen Machine Learning Cloud, INST 37/1057-1 FUGG. PB is a member of the Else Kröner Medical Scientist Kolleg “ClinbrAIn: Artificial Intelligence for Clinical Brain Research”. LK is supported by the Diabetes Center Berne. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting SM.
@@ Appendix A Toy example for disentanglement loss
Using a simple toy example, we showed that distance correlation can measure nonlinear dependencies between variables compared to two linear dependence measures. We sampled 1,000 points from two 2-dimensional subspaces $w_{1}=[w_{1,1},w_{1,2}]\in\mathbb{R}^{2}$ and $w_{2}=[w_{2,1},w_{2,2}]\in\mathbb{R}^{2}$ and created dependencies between them (Fig. A.1a, main diagonals in the pair plots). We defined an optimization problem, where points should be moved in order to minimize different dependence measures (Fig. A.1b-d). We optimized point coordinates with PyTorch (Paszke et al., 2019) and an Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.99,\epsilon=10^{-8}$) with a learning rate of 0.05 for 500 steps. When we optimized for the linear dependence measures Gaussian mutual information (GMI) and collapsed Gaussian mutual information (C-GMI), distance correlation as a nonlinear dependence measure only decreased for linear dependencies (Fig. A.1b-c, distance correlation values below points). GMI and C-GMI are two mutual information estimators that assume that the subspace vectors $w_{1}$ and $w_{2}$ follow a multivariate normal distribution (see A.1 for details). However, when we optimized for distance correlation, the dependence measure also dropped for nonlinear dependencies (Fig. A.1d, distance correlation values below points).
Figure A.1: Toy example for different dependence measures. We sampled 1,000 points from two 2-dimensional subspaces $w_{1}=[w_{1,1},w_{1,2}]\in\mathbb{R}^{2}$ and $w_{2}=[w_{2,1},w_{2,2}]\in\mathbb{R}^{2}$ and created dependencies between the main diagonals in the pair plots of a. We defined an optimization problem where we changed point coordinates to minimize various dependence measures. Below the points we show distance correlation as a nonlinear dependence measure. In panels b and c we show the displacement of points when minimizing the linear dependence measures GMI and C-GMI, respectively. In panel d we see that, compared to GMI and C-GMI, distance correlation can also effectively disentangle nonlinear dependencies between subspaces.
@@ A.1 Gaussian mutual information
$\text{GMI}(w_{1},w_{2})$ $=\frac{1}{2}\left(\log\left(\frac{\det\Sigma_{w_{1}}\det\Sigma_{w_{2}}}{\det\Sigma}\right)\right)$
@@
Approximating this measure over batches leads to numerically instabilities for covariance matrix estimation. Hence, we simplified the computation by taking the sum of the variables over their feature dimension $\hat{w_{1}}=\Sigma_{i=0}^{d_{1}}w_{1,i}$ and $\hat{w_{2}}=\Sigma_{i=0}^{d_{2}}w_{2,i}$. Since we collapse feature dimensions here, we name this measure collapsed Gaussian mutual information (C-GMI).
@@
Linear dependence measure.
When we assume $w_{1}$ and $w_{2}$ follow a multi-variate normal distribution, the GMI measure is only dependent on the covariance $\Sigma$ of the joint distribution $p(w_{1},w_{2})$
@@
Therefore, this measure only considers linear dependencies in the non-collapsed and collapsed form.
@@ Appendix E Failure cases
We qualitatively investigated the limitation of our generative model by checking for failure cases of a model trained for age-camera-style subspace disentanglement. We randomly sampled latent vectors from a standard normal distribution $z\sim\mathcal{N}(0,I)\in\mathcal{R}^{32}$ over different training epochs. We chose our best model in terms of validation loss. At the beginning of the training, when the generative model had not yet converged, the image features were cartoon-like (Fig. E.2a, epoch 9 and 17). In later stages of training, the fundus images got more realistic, however, we discovered some unrealistic features, such as images that were not perfectly circular (Fig. E.2b, first row, left edges), second optic discs (Fig. E.2b, first row, right center), or extreme colors and shapes (Fig. E.2b, second row). Sometimes images were generated with the non-black pixels in the background (Fig. E.2c) and few images even missed important anatomical structures such as the optic disc and the macula (Fig. E.2d). In some cases, we saw fundus images where the optic disc was on the wrong side (Fig. E.2e), and sometimes the images even looked very similar even though they were generated from different latent spaces (Fig. E.2e, first versus second row). We even discovered a few blurry images of poor quality (Fig. E.2f).
@@ Appendix F Different image resolutions
We trained generative models with age-camera-style disentanglement with different image resolutions and compared their performance with several metrics: disentanglement measure as the mean of subspace distance correlation values, age and camera encoding, both measured as the above chance level of accuracy, and the FID score as a metric of image quality. Here, we could see that our disentanglement measure was robust over different image resolutions (Fig. F.3, a), which was not surprising, since we used the same 32-dimensional latent space for each resolution. Age- and camera-encoding, on the other hand, improves for higher resolution (Fig. F.3, a). One possible explanation could be that it is easier to extract age and camera features from images with higher resolution. Finally, the FID score became worse (higher) with increasing resolution (Fig. F.3, a). Better (lower) FID scores tend to show greater sample diversity, but the higher resolution images showed qualitatively better image quality at higher resolution (Fig. F.3, b). Because we used the same 32-dimensional latent space for each image resolution, larger latent spaces may be required for better FID scores at higher image resolutions.
@@ Appendix G Multiple subspaces
To show that our model can handle combinations of multiple attributes, we trained generative models with subspaces for age, ethnicity, and camera encoding simultaneously. Therefore, compared to our previous experiments with 3 subspaces, we scaled up to 4 subspaces, resulting in 6 subspace combinations for distance correlation minimization (Eq. 4). Here we could see that despite training on 4 subspaces, our model still showed comparable performance: the kNN classifier performances on off-diagonals in our confusion matrix of label-subspace combinations for the models with disentanglement loss dropped similarly as in the experiments with only 3 subspaces (Fig. G.4, a versus b). In addition, the values of the distance correlation between the individual subspace combinations dropped from an average of 0.42 to 0.28, without major outliers (Fig. G.4, c).
Figure G.4: Disentanglement performance of a generative model with multiple subspaces. We trained generative models on retinal images, featuring four subspaces: $w_{\text{age}}$ encoding the age class $y_{\text{age}}$, $w_{\text{ethn}}$ encoding patient ethnicity $y_{\text{ethn}}$, $w_{\text{cam}}$ encoding the camera class $y_{\text{cam}}$, and $w_{\text{style}}$ serving as a patient-style subspace. In a and b, we report kNN classifier accuracy improvement over chance level accuracy. The baseline model in a involved training linear classifiers on the first three subspaces. In b, we additionally disentangled the subspaces using our distance correlation loss. Panel c compares the distance correlation measure between subspaces for the two model configurations.